# Construction of a Basis of Kähler differentials for generic plane branches

### v1

@@ 2.1. The set of exponents of an equisingularity class
So, up to a change of coordinates, we suppose that $\Gamma_{\circ}$ has a Puiseux expansion of the form (1). The calculation of the $\mathbb{C}[[x]]$-module of Kähler differentials on $\Gamma_{\circ}$ of $1$-forms preserving $E$ is simple if $\Gamma_{\circ}$ is smooth and transverse to $x=0$. Indeed, it is generated by $dx$, $dy$ or $xdy$ depending on whether $\{y=0\}\not\subset E$, $E=\{y=0\}$ or $E=\{xy=0\}$ respectively. Thus, we will assume from now on $n>1$.
@@ 2.2. The semigroup and the semi-module
The set $\Lambda_{\Gamma}^{E}$ is not a semigroup, but it is a so called semimodule: if $\ell\in\Lambda_{\Gamma}^{E}$ and $s\in\mathcal{S}(\Gamma)$, then obviously $\ell+s\in\Lambda_{\Gamma}^{E}$ (because $f\omega$ belongs to $\hat{\Omega}_{E}(\mathbb{C}^{2},0)$ if $\omega$ does). However, we do not need such generality. We shall only use the fact that if $\ell\in\Lambda_{\Gamma}^{E}$, then $n+\ell\in\Lambda_{\Gamma}^{E}$. If we write $[\ell]_{m}$ for the class modulo $m$ of $\ell$, then: if $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{n})$ is a family of $1$-forms such that $\nu_{\Gamma}(\Omega_{i})$ is minimum in $[\nu_{\Gamma}(\Omega_{i})]_{n}\cap\Lambda_{\Gamma}^{E}$ for $i=1,\ldots,n$ and $[\nu_{\Gamma}(\Omega_{i})]_{n}\neq[\nu_{\Gamma}(\Omega_{j})]_{n}$ for $i\neq j$, then certainly $\mathbf{\Omega}$ gives, by push-back to $\Gamma$, a basis of Kähler $1$-forms over $\Gamma$ and preserving $E$, i.e.
@@ 2.4. Approximate roots of $(\Gamma,E)$
In order to provide bases of Kähler differentials for generic curves in $\mathcal{C}$, we shall need to use both the elementary $1$-forms $dx$ and $dy$ and the approximate roots of $\Gamma$ where $\Gamma\in{\mathcal{C}}$.
Fix $11\leq j\leq g+$ 1. We define
@@
of $\Gamma$ up to (but not including) $\beta_{j}$. By convention, we denote $\beta_{g+1}=\infty$ and we obtain $\Gamma_{<\beta_{g+1}}=\Gamma$. By construction, $f_{j,\Gamma}\circ\Gamma_{<\beta_{j}}\equiv 0$ holds for all $\Gamma\in\mathcal{C}$ and $11\leq j\leq g+$ 1.
@@ 3. Case $E=\emptyset$
We divide our study into two cases: $E=\emptyset$, $E\neq\emptyset$. The former will in fact contain all the important definitions and arguments, whereas the latter will just consist in a generalization in which everything except the bootstrapping process carries over without modification.
@@ 3.1. Distinguished modules of $1$-forms
All the $1$-forms we shall construct will be of some level $l$, as the reader will see. Certainly, $dx$ and $dy$ are both of level 00.
@@ 3.2. Ordered families
At level 00, the first condition will be accomplished by considering the 1-forms $dx$ and $df_{1}$, that constitute a separated family. Then we describe a process in which $dx$ and $df_{1}$ appear alternatively and the approximate roots of $\Gamma$, with $\Gamma\in\mathcal{C}$, play a prominent role. We obtain $f_{1}=y$, since $\beta_{0}=\beta_{1}$ for the case $E=\emptyset$. Anyway, we have $f_{1}\neq y$ if $\beta_{0}<\beta_{1}$. So, in order to have uniform notations for the cases $E=\emptyset$ and $E\neq\emptyset$, we prefer to use $f_{1}$ in all cases.
@@ 3.3. The basic step — leading variables
Fix an equisingularity class ${\mathcal{C}}$ of germs of branches. Our construction essentially follows this process: assume we have found a terminal family (Definition 3.10) $\mathcal{T}_{l}=(\Omega_{1},\ldots,\Omega_{2{\nu}_{l}})$ in ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ of level $l<g$. Such family is separated but we do not delve now into its properties, which will be introduced in their definition. For now, a terminal family of level $l$ can be considered just as the product of our algorithm at the end of level $l$. We define $\mathcal{T}_{0}=\{dx,df_{1}\}$ as the terminal family of level 00. We are going to construct, iteratively, a terminal family $\mathcal{T}_{l+1}$ of level $l+1$ in ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$, with $2{\nu}_{l+1}$ elements. In the end, we shall have $\mathcal{T}_{g}$ with $2n$ elements whose first $n$ elements constitute a nice basis for $\mathcal{C}$. The construction of $\mathcal{T}_{l+1}$ requires several stages. The first one (stage 00) is elementary: we combine the $1$-forms in $\mathcal{T}_{l}$ with the approximate root $f_{l+1}$. More precisely, we define
@@
Fix a terminal family $\mathcal{T}_{l}=\{\Omega_{1},\ldots,\Omega_{2{\nu}_{l}}\}$ We say that
[…]
is a the initial family (of stage 00) of level $l+1$.
@@
Given $l\in{\mathbb{Z}}$ and $m\in{\mathbb{Z}}_{\geq 1}$, we denote by $[l]_{m}$ the class of $l$ modulo $m$. Moreover, given $S\subset{\mathbb{Z}}$, we define $[S]_{m}=\{[l]_{m}\ |\ \ l\in S\}$.
@@
Assume a family $\mathcal{G}_{l+1}^{s}=(\Omega_{1}^{s},\ldots,\Omega_{2{\nu}_{l+1}}^{s})$ is given, with $0\leq l<g$ and $s\geq 0$ (i.e. a family of level $l+1$ at some stage). Assume that $\mathcal{G}_{l+1}^{s}$ is non-separated. When this happens, we consider the pairs $j,k$ with $1\leq j<k\leq 2{\nu}_{l+1}$ such that $[\nu_{\mathcal{C}}(\Omega^{s}_{j})]_{n}=[\nu_{\mathcal{C}}(\Omega^{s}_{k})]_{n}$: whenever this equality holds, there is some $d\in\mathbb{Z}_{\geq 0}$ with $\nu_{\mathcal{C}}(\Omega^{s}_{k})=\nu_{\mathcal{C}}(\Omega_{j}^{s})+dn$. Consider the pull-backs
@@
Setting also $\Omega^{s+1}_{j}=\Omega^{s}_{j}$, for the other $1$-form, we obtain the $2{\nu}_{l+1}$ $1$-forms in ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ of stage $s+1$ (hence the exponent in $\Omega^{s+1}_{j}$), in the same level $l+1$.
@@
We have $\nu_{\mathcal{C}}(\Omega^{s+1}_{k})>\nu_{\mathcal{C}}(\Omega^{s}_{k})$ and $[\nu_{\mathcal{C}}(\Omega^{s+1}_{k})]_{n}\neq[\nu_{\mathcal{C}}(\Omega^{s}_{k})]_{n}$ as we shall see (Propositions 3.3 and 3.4).
@@
The algorithm described above to obtain a separated family $\mathcal{G}_{l+1}^{r}$ was called simple in section 1. One reason is that $\mathcal{G}_{l+1}^{0}$ is built from $\mathcal{T}_{l}$ and data related to the equisingularity class, namely the approximate root $f_{l+1}$. Another reason is that the algorithm consists in multiple applications of the elementary substitution (6). Indeed, the simplicity and expressiveness of Equation (6) will be essential to control the properties of our construction.
@@
  • If $\beta\not\in\left\{\beta_{1},\ldots,\beta_{g}\right\}$ (i.e. it is not a Puiseux characteristic exponent) and $\ell=\max\{l\geq 0:\beta_{l}\leq\beta\}$, then $m=0$ and $[k]_{e_{\ell}}=[0]_{e_{\ell}}$.
  • If $\beta=\beta_{\ell+1}$ (i.e. it is a Puiseux characteristic exponent), then $[k]_{e_{\ell}}=[m\beta]_{e_{\ell}}$.
@@
belong to $\mathbb{C}[a_{9},a_{10},\ldots][[t]]$, where we skip the subindexes $\Gamma$ in the second expression. In the former case, $t\frac{\partial(f_{1}\circ\Gamma(t))}{\partial t}$ has leading variable $a_{9}$ of degree 00 and we have $k=0$, $e_{0}=6$, $\tilde{c}=9$, $R\equiv 1$, $Q_{9}\equiv Q_{10}\equiv\ldots\equiv 0$, $c_{9}=1$, $c_{10}=10/9$, $c_{11}=11/9$ and so on.
@@
The nonzero polynomial $R(a_{\beta_{0}},\ldots,a_{{\mathfrak{p}}(\beta)})$ is the same for all the coefficients $C_{\beta}$ (this is what makes the definition of the $c-$sequence possible). Also, despite saying that $a_{\beta}$ is of “degree $m$”, the first element $C_{\beta}$ is
@@
But the terminology reflects the fact that $a_{\beta}$ is the maximal variable in the coefficient of the least order (the “leader”), and if $m>0$, then it subsequently appears explicitly always to the power $m$ in $C_{{\mathfrak{n}}^{r}(\beta)}$, for any $r>0$. If $\beta$ is not a Puiseux characteristic exponent, then $m=0$.
@@
This property, together with Remark 3.2, determine the $\mathcal{C}$-orders of all the $1$-forms appearing in the construction of a nice basis of $\mathcal{C}$. We obtain an algorithm to calculate such $\mathcal{C}$-orders without calculating the corresponding $1$-forms. This is the property of rich structure of the construction mentioned in section 1.
@@
Assume $\Omega\in\hat{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ (resp. $f\in\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[x,y]]$) belongs to $\Delta_{\beta_{l},\leq}^{m}$ for some $1\leq l\leq g$ and $20\leq m\leq n_{l}-$ 2. Then $\Omega\in\tilde{\Delta}_{\beta_{l},\leq}^{m}$ (resp. $f\in\tilde{\Delta}_{\beta_{l},\leq}^{m}$).
@@
By definition, we know that $[k]_{e_{l-1}}=[m\beta_{l}]_{e_{l-1}}$. Let $\Gamma$ be a generic curve with irreducible parametrization
@@
Assume, aiming at contradiction, that $Q_{\beta_{l}}\not\equiv 0$. This implies $C_{\beta_{l}}\neq 0$ for a generic curve as above. As the left hand side has order a multiple of $e_{l-1}$, we obtain $[0]_{e_{l-1}}=[k+\beta_{l}]_{e_{l-1}}$ and thus $(m+1)\beta_{l}$ is a multiple of $e_{l-1}$, which is impossible because $m+1$ is not a multiple of $n_{l}$. ∎
@@ 3.4. Leading variables of products
We first prove that product of several (say $a$) power series starting at a Puiseux characteristic exponent $\beta_{l}$ with degree 00 and $b$ power series starting at $\beta<\beta_{l}$ gives a power series with leading variable $\beta_{l}$ of degree $a-1$. We can also control the $c$-sequence of such product.
@@
[…]
@@
Finally, for products of powers of approximate roots and $1$-forms with the same leading variable with degree 00, we can also control the outcome, and the $c$-sequence. The next result is a consequence of Corollary 3.1, Lemma 3.2 and Proposition 3.1.
@@
All these results will be relevant for what will be stage 00 of each level $l+1$, in which we shall make use of both the $1$-forms constructed in levels up to $l$, and the approximate root $f_{l+1}$.

### v3

@@ 2.1. The set of exponents of an equisingularity class
So, up to a change of coordinates, we suppose that $\Gamma_{\circ}$ has a Puiseux expansion of the form (1). The calculation of the $\mathbb{C}[[x]]$-module of Kähler differentials on $\Gamma_{\circ}$ of $1$-forms preserving $E$ is simple if $\Gamma_{\circ}$ is smooth and transverse to $x=0$. Indeed, it is generated by $dx$, dydy or xdyxdy depending on whether $\{y=0\}\not\subset E$, $E=\{y=0\}$ or $E=\{xy=0\}$ respectively. Thus, we will assume from now on $n>1$.
@@ 2.2. The semigroup and the semi-module
The set $\Lambda_{\Gamma}^{E}$ is not a semigroup, but it is a so called semimodule: if $\ell\in\Lambda_{\Gamma}^{E}$ and $s\in\mathcal{S}(\Gamma)$, then obviously $\ell+s\in\Lambda_{\Gamma}^{E}$ (because $f\omega$ belongs to $\hat{\Omega}_{E}(\mathbb{C}^{2},0)$ if $\omega$ does). However, we do not need such generality. We shall only use the fact that if $\ell\in\Lambda_{\Gamma}^{E}$, then $n+\ell\in\Lambda_{\Gamma}^{E}$. If we write $\ell]_{m}$ for the class modulo $m$ of $\ell$, then: if $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{n})$ is a family of $1$-forms such that $\nu_{\Gamma}(\Omega_{i})$ is minimum in $\nu_{\Gamma}(\Omega_{i})]_{n}\cap\Lambda_{\Gamma}^{E}$ for $i=1,\ldots,n$ and $\nu_{\Gamma}(\Omega_{i})]_{n}\neq[\nu_{\Gamma}(\Omega_{j})]_{n}$ for $i\neq j$, then certainly $\mathbf{\Omega}$ gives, by push-back to $\Gamma$, a basis of Kähler $1$-forms over $\Gamma$ and preserving $E$, i.e.
@@ 2.4. Approximate roots of $(\Gamma,E)$
In order to provide bases of Kähler differentials for generic curves in $\mathcal{C}$, we shall need to use both the elementary $1$-forms dxdx and $dy$ and the approximate roots of $\Gamma$ where $\Gamma\in{\mathcal{C}}$.
Fix $1\leq j\leq g+1$. We define
@@
of $\Gamma$ up to (but not including) $\beta_{j}$. By convention, we denote $\beta_{g+1}=\infty$ and we obtain $\Gamma_{<\beta_{g+1}}=\Gamma$. By construction, $f_{j,\Gamma}\circ\Gamma_{<\beta_{j}}\equiv 0$ holds for all $\Gamma\in\mathcal{C}$ and $1\leq j\leq g+1$.
@@ 3. Case $E=\emptyset$
We divide our study into two cases: $\emptyset, EE\neq\emptyset$. The former will in fact contain all the important definitions and arguments, whereas the latter will just consist in a generalization in which everything except the bootstrapping process carries over without modification.
@@ 3.1. Distinguished modules of $1$-forms
All the $1$-forms we shall construct will be of some level $l$, as the reader will see. Certainly, $dx$ and $dy$ are both of level 0.
@@ 3.2. Ordered families
At level 0, the first condition will be accomplished by considering the 1-forms $dx$ and $df_{1}$, that constitute a separated family. Then we describe a process in which $dx$ and $df_{1}$ appear alternatively and the approximate roots of $\Gamma$, with $\Gamma\in\mathcal{C}$, play a prominent role. We obtain $f_{1}=y$, since $\beta_{0}=\beta_{1}$ for the case $E=\emptyset$. Anyway, we have $f_{1}\neq y$ if $\beta_{0}<\beta_{1}$. So, in order to have uniform notations for the cases $E=\emptyset$ and $E\neq\emptyset$, we prefer to use $f_{1}$ in all cases.
@@ 3.3. The basic step — leading variables
Fix an equisingularity class $\mathcal{C}$} of germs of branches. Our construction essentially follows this process: assume we have found a terminal family (Definition 3.10) $\mathcal{T}_{l}=(\Omega_{1},\ldots,\Omega_{2{\nu}_{l}})$ in ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ of level $l<g$. Such family is separated but we do not delve now into its properties, which will be introduced in their definition. For now, a terminal family of level $l$ can be considered just as the product of our algorithm at the end of level $l$. We define $\mathcal{T}_{0}=(dx,df_{1})$ as the terminal family of level 0. We are going to construct, iteratively, a terminal family $\mathcal{T}_{l+1}$ of level $l+1$ in ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$, with $2{\nu}_{l+1}$ elements. In the end, we shall have $\mathcal{T}_{g}$ with $2n$ elements whose first $n$ elements constitute a nice basis for $\mathcal{C}$. The construction of $\mathcal{T}_{l+1}$ requires several stages. The first one (stage 0) is elementary: we combine the $1$-forms in $\mathcal{T}_{l}$ with the approximate root $f_{l+1}$. More precisely, we define
@@
Fix a terminal family $\mathcal{T}_{l}=(\Omega_{1},\ldots,\Omega_{2{\nu}_{l}})$ . We say that
$\mathcal{G}_{l+1}^{0}=(\Omega^{0}_{2a\nu_{l}+j})_{0\leq a<n_{l+1},\ 1\leq j\leq 2{\nu}_{l}}=(\Omega^{0}_{k})_{1\leq k\leq 2{\nu}_{l+1}}$
is a the initial family (of stage 0) of level $l+1$.
@@
Given $l\in{\mathbb{Z}}$ and $m\in{\mathbb{Z}}_{\geq 1}$, we denote by $l]_{m}$ the class of $l$ modulo $m$. Moreover, given $S\subset{\mathbb{Z}}$, we define $S]_{m}=\{[l]_{m}\ |\ \ l\in S\}$.
@@
Assume a family $\mathcal{G}_{l+1}^{s}=(\Omega_{1}^{s},\ldots,\Omega_{2{\nu}_{l+1}}^{s})$ is given, with $0\leq l<g$ and $s\geq 0$ (i.e. a family of level $l+1$ at some stage). Assume that $\mathcal{G}_{l+1}^{s}$ is non-separated. When this happens, we consider the pairs $j,k$ with $1\leq j<k\leq 2{\nu}_{l+1}$ such that $\nu_{\mathcal{C}}(\Omega^{s}_{j})]_{n}=[\nu_{\mathcal{C}}(\Omega^{s}_{k})]_{n}$: whenever this equality holds, there is some $d\in\mathbb{Z}_{\geq 0}$ with $\nu_{\mathcal{C}}(\Omega^{s}_{k})=\nu_{\mathcal{C}}(\Omega_{j}^{s})+dn$. Consider the pull-backs
@@
Setting also $\Omega^{s+1}_{j}=\Omega^{s}_{j}$, for the other $1$-form, we obtain the $2{\nu}_{l+1}$ 11-forms in ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ of stage $s+1$ (hence the exponent in $\Omega^{s+1}_{j}$), in the same level $l+1$.
@@
We have $\nu_{\mathcal{C}}(\Omega^{s+1}_{k})>\nu_{\mathcal{C}}(\Omega^{s}_{k})$ and $\nu_{\mathcal{C}}(\Omega^{s+1}_{k})]_{n}\neq[\nu_{\mathcal{C}}(\Omega^{s}_{k})]_{n}$ as we shall see (Propositions 3.3 and 3.4).
@@
The algorithm described above to obtain a separated family $\mathcal{G}_{l+1}^{r}$ was described as simple in section 1. One reason is that $\mathcal{G}_{l+1}^{0}$ is built from $\mathcal{T}_{l}$ and data related to the equisingularity class, namely the approximate root $f_{l+1}$. Another reason is that the algorithm consists in multiple applications of the elementary substitution (6). Indeed, the simplicity and expressiveness of Equation (6) will be essential to control the properties of our construction.
@@
  • If $\beta\not\in\left\{\beta_{1},\ldots,\beta_{g}\right\}$ (i.e. it is not a Puiseux characteristic exponent) and $\ell=\max\{l\geq 0:\beta_{l}\leq\beta\}$, then $m=0$ and $k]_{e_{\ell}}=[0]_{e_{\ell}}$.
  • If $\beta=\beta_{\ell+1}$ (i.e. it is a Puiseux characteristic exponent), then $k]_{e_{\ell}}=[m\beta]_{e_{\ell}}$.
@@
belong to $\mathbb{C}[a_{9},a_{10},\ldots][[t]]$, where we skip the subindexes $\Gamma$ in the second expression. In the former case, $t\frac{\partial(f_{1}\circ\Gamma(t))}{\partial t}$ has leading variable $a_{9}$ of degree 0 and we have $k=0$, $e_{0}=6$, $\tilde{c}=9$, $R\equiv 1$, $Q_{9}\equiv Q_{10}\equiv\ldots\equiv 0$, $c_{9}=1$, $c_{10}=10/9$, $c_{11}=11/9$ and so on.
@@
The nonzero polynomial $R(a_{\beta_{0}},\ldots,a_{{\mathfrak{p}}(\beta)})$ is the same for all the coefficients $C_{\beta}$ (this is what makes the definition of the $c$-sequence possible). Also, despite saying that $a_{\beta}$ is of “degree mm”, the first element $C_{\beta}$ is
@@
But the terminology reflects the fact that $a_{\beta}$ is the variable of maximal index in the coefficient of the least order (the “leader”), and if $m>0$, then it subsequently appears explicitly always to the power $m$ in $C_{{\mathfrak{n}}^{r}(\beta)}$, for any $r>0$. If $\beta$ is not a Puiseux characteristic exponent, then $m=0$.
@@
This property, together with Remark 3.2, determine the $\mathcal{C}$-orders of all the $1$-forms appearing in the construction of a nice basis of $\mathcal{C}$. We obtain Algorithm 1 that calculates such $\mathcal{C}$-orders without calculating the corresponding $1$-forms. This is the property of rich structure of the construction mentioned in section 1.
@@
Assume $\Omega\in\hat{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ (resp. $f\in\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[x,y]]$) belongs to $\Delta_{\beta_{l},\leq}^{m}$ for some $1\leq l\leq g$ and $0\leq m\leq n_{l}-2$. Then $\Omega\in\tilde{\Delta}_{\beta_{l},\leq}^{m}$ (resp. $f\in\tilde{\Delta}_{\beta_{l},\leq}^{m}$).
@@
By definition, we know that $k]_{e_{l-1}}=[m\beta_{l}]_{e_{l-1}}$. Let $\Gamma$ be a generic curve with irreducible parametrization
@@
Assume, aiming at contradiction, that $Q_{\beta_{l}}\not\equiv 0$. This implies $C_{\beta_{l}}\neq 0$ for a generic curve as above. As the left hand side has order a multiple of $e_{l-1}$, we obtain $0]_{e_{l-1}}=[k+\beta_{l}]_{e_{l-1}}$ and thus $(m+1)\beta_{l}$ is a multiple of $e_{l-1}$, which is impossible because $m+1$ is not a multiple of $n_{l}$. ∎
@@ 3.4. Leading variables of products
We first prove that product of several (say $a$) power series starting at a Puiseux characteristic exponent $\beta_{l}$ with degree 0 and $b$ power series starting at $\beta<\beta_{l}$ gives a power series with leading variable $\beta_{l}$ of degree a1a-1. We can also control the $c$-sequence of such product.
@@
$g_{i}(t)\in\bigcup_{\beta^{\prime}<\beta_{l}}\Delta^{0}_{\beta^{\prime}}\ \mathrm{for}\ 1\leq i\leq\nu_{l-1}-1.$
@@
Finally, for products of powers of approximate roots and $1$-forms with the same leading variable with degree 0, we can also control the outcome, and the $c$-sequence. The next result is a consequence of Corollary 3.1, Lemma 3.2 and Proposition 3.1.
@@
All these results will be relevant for what will be stage 0 of each level $l+1$, in which we shall make use of both the $1$-forms constructed in levels up to $l$, and the approximate root $f_{l+1}$.